# Looking Beyond Visible Cues: Implicit Video Question Answering via Dual-Clue Reasoning

### v1

@@ I.2 Details of the PSAV Dataset
The “Persuasion Strategies in Advertisement Videos” dataset [37] (PSAV) contributes an important task on the advertisement story understanding task, namely persuasion strategy identification. For this task, the authors collect 1,002 video advertisements from popular brands available on the web publicly and utilize the persuasion strategy labels defined by [38]. The 12 strategies target persuasion strategy set are: Social Identity, Concreteness, Anchoring and Comparison, Overcoming Reactance, Reciprocity, Foot-in-the-Door, Authority, Social Impact, Anthropomorphism, Scarcity, Social Proof, and Unclear.
I.3 Details of the SUTD-Traffic Datasets
@@
The SUTD-TrafficQA [73] dataset is a large-scale collection of 10,080 videos from diverse real-world traffic scenarios, sourced both online (YouTube, LiveLeak, Twitter, Bili-bili) and offline (handheld cameras and car-mounted recorders) to ensure global diversity. It contains 62,535 human-generated question-answer pairs across six complex reasoning tasks: basic event understanding, event forecasting, reverse reasoning, counterfactual inference, introspection, and attribution. These tasks focus on causal inference, event understanding, and multi-level logical reasoning.
Appendix J Limitations and Future Work

### v2

@@ H.2 Details of the PSAV Dataset
The “Persuasion Strategies in Advertisement Videos” dataset [adsdataset] (PSAV) contributes an important task on the advertisement story understanding task, namely persuasion strategy identification. For this task, the authors collect 1,002 video advertisements from popular brands available on the web publicly and utilize the persuasion strategy labels defined by [Persuasion]. The 12 strategies target persuasion strategy set are: Social Identity, Concreteness, Anchoring and Comparison, Overcoming Reactance, Reciprocity, Foot-in-the-Door, Authority, Social Impact, Anthropomorphism, Scarcity, Social Proof, and Unclear.
H.3 Details of the SUTD-Traffic Datasets
@@
The SUTD-TrafficQA [sutd] dataset is a large-scale collection of 10,080 videos from diverse real-world traffic scenarios, sourced both online (YouTube, LiveLeak, Twitter, Bili-bili) and offline (handheld cameras and car-mounted recorders) to ensure global diversity. It contains 62,535 human-generated question-answer pairs across six complex reasoning tasks: basic event understanding, event forecasting, reverse reasoning, counterfactual inference, introspection, and attribution. These tasks focus on causal inference, event understanding, and multi-level logical reasoning.
Appendix I Limitations and Future Work